# VariFace: Fair and Diverse Synthetic Dataset Generation for Face Recognition

### v1

@@ 4 Experiments
SynFace [53] 0.5M ($10K\times 50$) 0.8407 0.6337 0.6355 0.5910 0.6937 0.6789 -0.2681
IDNet* [32] 0.5M ($10K\times 50$) 0.9258 0.7540 0.7425 0.6388 0.7990 0.7913 -0.1557
ExFaceGAN* [9] 0.5M ($10K\times 50$) 0.9350 0.7384 0.7160 0.7892 0.8298 0.8017 -0.1453
DigiFace* [2] 0.5M ($10K\times 50$) 0.9540 0.8740 0.7887 0.7697 0.7862 0.8345 -0.1125
VIGFace* [31] 0.5M ($10K\times 50$) 0.9660 0.8666 0.7503 0.8250 0.8342 0.8484 -0.0986
IDiff-Face* [8] 0.5M ($10K\times 50$) 0.9800 0.8547 0.8045 0.8643 0.9065 0.8820 -0.0650
DCFace* [30] 0.5M ($10K\times 50$) 0.9855 0.8533 0.8262 0.8970 0.9160 0.8956 -0.0514
ID3* [36] 0.5M ($10K\times 50$) 0.9768 0.8684 0.8277 0.9100 0.9073 0.8980 -0.0490
Arc2Face* [49] 0.5M ($10K\times 50$) 0.9881 0.9187 0.8516 0.9018 0.9263 0.9173 -0.0297
Vec2Face* [72] 0.5M ($10K\times 50$) 0.9887 0.8897 0.8547 0.9312 0.9357 0.9200 -0.0270
@@
4.1 Implementation details
For the diffusion model, we adapted the Hourglass Diffusion Transformer (HDiT) [12] architecture to handle multiple conditions (see Fig. S7). We trained the stage 1 and 2 diffusion models for 700K and 3M iterations, which took $\approx$19 hours and 3.5 days, respectively, using 4 NVIDIA A100 GPUs. During inference, we applied the DPM-Solver++(3M) SDE sampler [40, 12] with 50 sampling steps. To create demographic labels, we used the pretrained ViT-L-14 MetaCLIP model [74]. For FR consistency, we used an IResNet-100 model trained on the CASIA-WebFace dataset to extract embeddings and set $K=50$. Regarding filtering settings, we empirically set the quality threshold in stage 1 at $0.7$ and the cosine similarity threshold at $0.3$ in stage 2 [72]. For the stage 1 generation, we sampled a balanced set of individuals across races and genders. For the stage 2 generation, we uniformly sampled age values $A\sim\mathcal{U}(0,1)$ and divergence scores $DS\sim\mathcal{U}(0.5,0.8)$ (see Table. S2).
Within the synthetic generation pipeline, where performance is the main concern, we used an IResNet-100 model, while for evaluation, we used the smaller IResNet-50 architecture. For both models, we used the ArcFace loss function [14] and trained for 40 epochs, with an initial learning rate of 0.1, which is reduced by a factor of 10 at epochs 24, 30, and 36. Moreover, we applied the following data augmentations: horizontal flip, sharpness, contrast, equalization, and random erasing. More detailed information on the diffusion and FR model settings can be found in the Suppl. Sec. 11 and 12, respectively.
4.2 Datasets
For training, we used the CASIA-WebFace dataset [75], a publicly available face dataset containing 490,414 images from 10,575 individuals. For evaluation, we selected six common face verification datasets: LFW [22], CFP-FP [59], CPLFW [76], AgeDB [46], CALFW [77] and RFW dataset [70]. These datasets were designed to evaluate specific aspects of FR performance, including pose variation (CFP-FP, CPLFW), large age differences (AgeDB, CALFW), and race (RFW). Following previous methods [30, 31, 72], we group the LFW, CFP-FP, CPLFW, AgeDB, and CALFW datasets and refer to these as the ‘Standard Benchmark’, and we report the performance difference to a baseline model trained on CASIA-WebFace (Real Gap).
4.3 Constrained Face Recognition Results
@@
SynFace [53] 0.5643 0.6355 0.6647 0.6457 0.6276 -0.2710
DigiFace [2] 0.5952 0.6408 0.6750 0.6427 0.6384 -0.2602
DCFace [30] 0.7742 0.8122 0.8917 0.8460 0.8310 -0.0676
Vec2Face [72] 0.8415 0.8535 0.9028 0.8750 0.8682 -0.0304
@@ 4.4 Unconstrained Face Recognition Results
SynFace [53] 1.0M ($10K\times 100$) 0.8580 0.6473 0.6395 0.5765 0.6987 0.6840 -0.2630
DigiFace* [2] 1.2M ($10K\times 72+100K\times 5$) 0.9582 0.8877 0.8162 0.7972 0.8070 0.8532 -0.0938
DCFace* [30] 1.2M ($20K\times 50+40K\times 5$) 0.9858 0.8861 0.8507 0.9097 0.9282 0.9121 -0.0349
VIGFace* [31] 4.2M ($60\times 50+60\times 20$) 0.9913 0.9187 0.8483 0.9463 0.9338 0.9277 -0.0193
Arc2Face* [49] 1.2M ($20K\times 50+40K\times 5$) 0.9892 0.9458 0.8645 0.9245 0.9333 0.9314 -0.0156
Vec2Face* [72] 15M ($300K\times 50$) 0.9930 0.9154 0.8770 0.9445 0.9458 0.9352 -0.0118
@@
SynFace [53] 1.0M ($10K\times 100$) 0.5873 0.6617 0.6758 0.6502 0.6438 -0.2548
DigiFace [2] 1.2M ($10K\times 72+100K\times 5$) 0.6145 0.6648 0.6910 0.6603 0.6577 -0.2409
DCFace [30] 1.2M ($20K\times 50+40K\times 5$) 0.8157 0.8345 0.9083 0.8765 0.8588 -0.0398
Vec2Face [72] 1.0M ($20K\times 50$) 0.8763 0.8695 0.9185 0.8962 0.8901 -0.0085
@@
[…]
4.5 Synthetic dataset characteristics
Intraclass diversity. Histogram plots of DS for different synthetic datasets are shown in Fig. 4. Low cosine similarity values generally correspond to label noise, representing either mislabeled data for real datasets or lack of ID preservation for synthetic datasets (see Suppl. Sec. 8). Among the synthetic datasets, SynFace displays many images with low cosine similarity values due to the use of ID mixing [53]. In contrast, other synthetic methods contain mainly high cosine similarity values, corresponding to low intraclass diversity. VariFace controls the DS to enable the generation of diverse image variation while preserving ID.
4.6 Ablation study
To assess the effect of each conditioning signal used in VariFace, we evaluate the FR performance trained with VariFace using different combinations of conditions (Table 5). By removing the race and gender conditioning in stage 1, the performance on the Standard Benchmark remains comparable ($0.9405\rightarrow 0.9411$), but there is a considerable decrease in RFW performance ($0.8978\rightarrow 0.8820$), especially with minority races (Asian: $0.8733\rightarrow 0.8365$, Indian: $0.8988\rightarrow 0.8683$, African: $0.8895\rightarrow 0.8662$). Both age and DS conditioning improve face verification accuracy across all datasets, with age conditioning improving performance on the AgeDB and CALFW datasets, and DS conditioning improving performance on the CFP-FP and CPLFW datasets. Overall, the best performance is achieved using all three conditions, providing fairness across races and robustness to significant pose and age variation.
@@ 5 Conclusions
In this paper, we propose VariFace, a two-stage, diffusion-based pipeline for generating fair and diverse synthetic face datasets for training models. We introduce Face Recognition Consistency to refine demographic labels, Face Vendi Score Guidance to improve interclass diversity, and Divergence Score Conditioning to balance the ID preservation-intraclass diversity trade-off. When controlling for the dataset size and the number of images per ID, VariFace consistently outperforms previous synthetic dataset methods across six evaluation datasets and achieves FR performance comparable to real data (Table 1, 2). Furthermore, by scaling synthetic dataset size, VariFace outperforms the real dataset, as well as previous synthetic methods at all dataset sizes (Table 3, 4). Importantly, VariFace addresses fairness concerns, achieving better representation of minority demographic classes demonstrated both through qualitative visualizations (Fig. 4) and quantitative evaluation (Table. 2).
@@
With a gap between the FR performances from training on synthetic and real datasets, previous methods have emphasized the benefit of using synthetic data to augment real datasets [53, 2, 31]. In contrast, we demonstrate for the first time that state-of-the-art FR performance can be achieved when training only on synthetic data. Therefore, our results establish synthetic face datasets as a viable solution for achieving accurate and fair FR performance.
@@
Acknowledgements. We are grateful to Prof. Satoshi Hara of the University of Electro-Communications for reviewing this manuscript.
Supplementary Material
6 CLIP Demographic labeling
6.1 Framework
@@
We adapt the CLIP-IQA framework [68] to obtain demographic labels without relying on supervised models (Fig. S1). Unlike CLIP-IQA, we input both the image and its horizontally flipped version into the CLIP image encoder to obtain their corresponding image embeddings. We use the horizontally flipped image as a form of test-time augmentation to improve the robustness of the predictions. For the text encoder, we pass a set of prompts with the structure “A photo of a * face” where * is replaced with [“Male”, “Female”], [“Young”, “Old”], [“Caucasian”, “Asian”, “Indian”, “African”] for gender, age and race labeling, respectively. Given a set of text embeddings $T=\{t_{1},t_{2},...,t_{n}\}$ and the embedding of an image and its horizontally flipped version $X=\{x_{1},x_{2}\}$, first the cosine similarity is computed:
@@
6.2 Evaluation
To evaluate the benefit of using CLIP compared to supervised models, we compare CLIP with DeepFace [60], a library that contains supervised models for race prediction of face images. Furthermore, we evaluate the benefit of including Face Recognition Consistency (FRC) with CLIP predictions.
For evaluation, we use the RFW dataset, which contains 48,000 images evenly divided across the following races: [“Caucasian”, “Asian”, “Indian”, “African”]. DeepFace includes “Latino Hispanic” and “Middle Eastern” as predicted classes, and we relabel these as “Indian” predictions to conform to the RFW race categories. For FRC, we set $K=10$ to account for the small dataset size. The accuracy of the predictions using DeepFace, CLIP, and CLIP-FRC for each race are shown in Fig. S2.

### v2

@@ 4 Experiments
4.1 Implementation details
For the diffusion model, we adapted the Hourglass Diffusion Transformer (HDiT) (Crowson et al., 2024) architecture to handle multiple conditions (see Fig. S7). We trained the stage 1 and 2 diffusion models for 700K and 3M iterations, which took $\approx$19 hours and 3.5 days, respectively, using 4 NVIDIA A100 GPUs. During inference, we applied the DPM-Solver++(3M) SDE sampler (Lu et al., 2022; Crowson et al., 2024) with 50 sampling steps. To create demographic labels, we used the pretrained ViT-L-14 MetaCLIP model (Xu et al., 2023). For FR consistency, we used an IResNet-100 model trained on the CASIA-WebFace dataset to extract embeddings and set $K=50$. Regarding filtering settings, we empirically set the quality threshold in stage 1 at $0.7$ and the cosine similarity threshold at $0.3$ in stage 2 (Wu et al., 2024a). For the stage 1 generation, we sampled a balanced set of individuals across races and genders. For the stage 2 generation, we uniformly sampled age values $A\sim\mathcal{U}(0,1)$ and divergence scores $DS\sim\mathcal{U}(0.5,0.8)$ (see Table S2).
Within the synthetic generation pipeline, where performance is the main concern, we used an IResNet-100 model, while for evaluation, we used the smaller IResNet-50 architecture. For both models, we used the ArcFace loss function (Deng et al., 2019) and trained for 40 epochs, with an initial learning rate of 0.1, which is reduced by a factor of 10 at epochs 24, 30, and 36. Moreover, we applied the following data augmentations: horizontal flip, sharpness, contrast, equalization, and random erasing. More detailed information on the diffusion and FR model settings can be found in the Suppl. Sec. F and G, respectively.
4.2 Datasets
For training, we used the CASIA-WebFace dataset (Yi et al., 2014), a publicly available face dataset containing 490,414 images from 10,575 individuals. For evaluation, we selected six common face verification datasets: LFW (Huang et al., 2008), CFP-FP (Sengupta et al., 2016), CPLFW (Zheng & Deng, 2018), AgeDB (Moschoglou et al., 2017), CALFW (Zheng et al., 2017) and RFW dataset (Wang et al., 2019). These datasets were designed to evaluate specific aspects of FR performance, including pose variation (CFP-FP, CPLFW), large age differences (AgeDB, CALFW), and race (RFW). Following previous methods (Kim et al., 2023; 2024; Wu et al., 2024a), we group the LFW, CFP-FP, CPLFW, AgeDB, and CALFW datasets and refer to these as the ‘Standard Benchmark’, and we report the performance difference to a baseline model trained on CASIA-WebFace (Real Gap).
4.3 Constrained Face Recognition Results
@@
SynFace (Qiu et al., 2021) 0.5M ($10K\times 50$) 0.8407 0.6337 0.6355 0.5910 0.6937 0.6789 -0.2681
IDNet* (Kolf et al., 2023) 0.5M ($10K\times 50$) 0.9258 0.7540 0.7425 0.6388 0.7990 0.7913 -0.1557
ExFaceGAN* (Boutros et al., 2023b) 0.5M ($10K\times 50$) 0.9350 0.7384 0.7160 0.7892 0.8298 0.8017 -0.1453
DigiFace* (Bae et al., 2023) 0.5M ($10K\times 50$) 0.9540 0.8740 0.7887 0.7697 0.7862 0.8345 -0.1125
VIGFace* (Kim et al., 2024) 0.5M ($10K\times 50$) 0.9660 0.8666 0.7503 0.8250 0.8342 0.8484 -0.0986
IDiff-Face* (Boutros et al., 2023a) 0.5M ($10K\times 50$) 0.9800 0.8547 0.8045 0.8643 0.9065 0.8820 -0.0650
DCFace* (Kim et al., 2023) 0.5M ($10K\times 50$) 0.9855 0.8533 0.8262 0.8970 0.9160 0.8956 -0.0514
ID3* (Li et al., 2024) 0.5M ($10K\times 50$) 0.9768 0.8684 0.8277 0.9100 0.9073 0.8980 -0.0490
Arc2Face* (Papantoniou et al., 2024) 0.5M ($10K\times 50$) 0.9881 0.9187 0.8516 0.9018 0.9263 0.9173 -0.0297
Vec2Face* (Wu et al., 2024a) 0.5M ($10K\times 50$) 0.9887 0.8897 0.8547 0.9312 0.9357 0.9200 -0.0270
@@
SynFace (Qiu et al., 2021) 0.5643 0.6355 0.6647 0.6457 0.6276 -0.2710
DigiFace (Bae et al., 2023) 0.5952 0.6408 0.6750 0.6427 0.6384 -0.2602
DCFace (Kim et al., 2023) 0.7742 0.8122 0.8917 0.8460 0.8310 -0.0676
Vec2Face (Wu et al., 2024a) 0.8415 0.8535 0.9028 0.8750 0.8682 -0.0304
@@ 4.4 Unconstrained Face Recognition Results
SynFace (Qiu et al., 2021) 1.0M ($10K\times 100$) 0.8580 0.6473 0.6395 0.5765 0.6987 0.6840 -0.2630
DigiFace* (Bae et al., 2023) 1.2M ($10K\times 72+100K\times 5$) 0.9582 0.8877 0.8162 0.7972 0.8070 0.8532 -0.0938
DCFace* (Kim et al., 2023) 1.2M ($20K\times 50+40K\times 5$) 0.9858 0.8861 0.8507 0.9097 0.9282 0.9121 -0.0349
VIGFace* (Kim et al., 2024) 4.2M ($60\times 50+60\times 20$) 0.9913 0.9187 0.8483 0.9463 0.9338 0.9277 -0.0193
Arc2Face* (Papantoniou et al., 2024) 1.2M ($20K\times 50+40K\times 5$) 0.9892 0.9458 0.8645 0.9245 0.9333 0.9314 -0.0156
Vec2Face* (Wu et al., 2024a) 15M ($300K\times 50$) 0.9930 0.9154 0.8770 0.9445 0.9458 0.9352 -0.0118
@@
SynFace (Qiu et al., 2021) 1.0M ($10K\times 100$) 0.5873 0.6617 0.6758 0.6502 0.6438 -0.2548
DigiFace (Bae et al., 2023) 1.2M ($10K\times 72+100K\times 5$) 0.6145 0.6648 0.6910 0.6603 0.6577 -0.2409
DCFace (Kim et al., 2023) 1.2M ($20K\times 50+40K\times 5$) 0.8157 0.8345 0.9083 0.8765 0.8588 -0.0398
Vec2Face (Wu et al., 2024a) 1.0M ($20K\times 50$) 0.8763 0.8695 0.9185 0.8962 0.8901 -0.0085
@@ 4.5 Synthetic dataset characteristics
Intraclass diversity. Histogram plots of DS for different synthetic datasets are shown in Fig. 4. Low cosine similarity values generally correspond to label noise, representing either mislabeled data for real datasets or lack of ID preservation for synthetic datasets (see Suppl. Sec. C). Among the synthetic datasets, SynFace displays many images with low cosine similarity values due to the use of ID mixing (Qiu et al., 2021). In contrast, other synthetic methods contain mainly high cosine similarity values, corresponding to low intraclass diversity. VariFace controls the DS to enable the generation of diverse image variation while preserving ID.
4.6 Dataset generation time
In addition to performance improvements, VariFace also takes less time to generate face datasets than previous methods (Table 5).
Method ()($\downarrow$)( ↓ ) Dataset Generation Time (hours)
DCFace (Kim et al., 2023) 20
Vec2Face (Wu et al., 2024a) 36
VariFace (ours) 12
Table 5: Dataset generation time. 500K images generated using second stage generative model on a single NVIDIA A100 GPU.
The entire pipeline to generate 500K images using VariFace, including ID generation and filtering, took $\approx$15 hours. There are three main reasons for Variface’s efficiency. Firstly, while FVSG adds inference overhead to stage 1 generation, this cost is minimal because face datasets contain significantly fewer IDs than images. Therefore, the stage 1 generation is fast, taking $\approx$15 minutes to generate 10,000 IDs without FVSG, or $\approx$40 minutes with FVSG. The main bottleneck is in the second stage, where multiple images of the same ID are generated. To minimize inference costs in this stage, our proposed DSC provides a lightweight solution to generate diversity without relying on additional gradient computations through auxiliary models. Finally, we adopt an efficient diffusion architecture and fast sampling method that reduces inference time across both stages.
4.7 Ablation study
@@
Conditioning signal. To assess the effect of each conditioning signal used in VariFace, we evaluate the FR performance trained with VariFace using different combinations of conditions (Table 6). By removing the race and gender conditioning in stage 1, the performance on the Standard Benchmark remains comparable ($0.9405\rightarrow 0.9411$), but there is a considerable decrease in RFW performance ($0.8978\rightarrow 0.8820$), especially with minority races (Asian: $0.8733\rightarrow 0.8365$, Indian: $0.8988\rightarrow 0.8683$, African: $0.8895\rightarrow 0.8662$). Both age and DS conditioning improve face verification accuracy across all datasets, with age conditioning improving performance on the AgeDB and CALFW datasets, and DS conditioning improving performance on the CFP-FP and CPLFW datasets. Overall, the best performance is achieved using all three conditions, providing fairness across races and robustness to significant pose and age variation.
Face Vendi Score Guidance. The purpose of FVSG is to improve the interclass diversity (see Suppl Sec. B). Previous methods to improve interclass diversity relied on ID filtering (Kim et al., 2023; Wu et al., 2024a), where FR embeddings from synthesized identities with a cosine similarity above a predefined threshold are removed. Here, we demonstrate that FVSG improves face verification accuracy and is a better alternative to ID filtering (Table 7), with the most improvement observed on the RFW dataset ($0.8935\rightarrow 0.8978$). In contrast, we observe that ID filtering was associated with a decrease in performance across datasets ($0.8935\rightarrow 0.8927$). While ID filtering and FVSG are designed for the same purpose, ID filtering does not encourage sampling more diverse IDs and is sensitive to the FR model and threshold used. Moreover, the removal of similar IDs using ID filtering risks removing difficult-to-classify but distinct IDs misclassified by the FR model, which may be most informative for training (Swayamdipta et al., 2020). In contrast, FVSG ensures all demographic groups are well represented by directly optimizing the distribution of generated identities. Furthermore, FVSG is less dependent on the FR model because the model is used as a guidance signal without thresholding and does not risk removing informative cases.
FVSG S1F LFW CFP-FP CPLFW AgeDB CALFW Average African Asian Caucasian Indian Average
Q 0.9938 0.9460 0.8882 0.9438 0.9305 0.9405 0.8895 0.8733 0.9295 0.8988 0.8978
Q 0.9920 0.9433 0.8917 0.9412 0.9313 0.9399 0.8783 0.8780 0.9260 0.8917 0.8935
Q+ID 0.9923 0.9451 0.8897 0.9388 0.9303 0.9393 0.8850 0.8695 0.9233 0.8928 0.8927
Table 7: FVSG ablation study. FVSG=Face Vendi Score Guidance. S1F=Stage 1 filtering. Q=quality. ID=identity. All datasets contain 0.5M images (50 images/ID).
Filtering. Filtering is widely used in synthetic face generation Kim et al. (2023); Wu et al. (2024a), although the importance of filtering has not been well studied. Similar to Wu et al. (2024a), we noticed that filtering only affected a small proportion of the generated images, and here we further investigated the impact of filtering on face verification accuracy (Table 8). While the best overall performance is observed with filtering, we observe that filtering is not necessary for VariFace to achieve state-of-the-art results. Even without any filtering, VariFace still considerably outperforms previous SOTA methods ($0.9386>0.9200$), highlighting the robustness of the conditional generation models used in VariFace.
S1F S2F LFW CFP-FP CPLFW AgeDB CALFW Average African Asian Caucasian Indian Average
0.9937 0.9410 0.8892 0.9382 0.9308 0.9386 0.8867 0.8785 0.9247 0.8962 0.8965
0.9933 0.9444 0.8923 0.9413 0.9325 0.9408 0.8880 0.8722 0.9272 0.8918 0.8948
0.9908 0.9464 0.8912 0.9353 0.9293 0.9386 0.8828 0.8713 0.9212 0.8912 0.8916
Table 8: Filtering ablation study. SB=Standard Benchmark. S1F=Stage 1 filtering. S2F=Stage 2 filtering. All datasets contain 0.5M images (50 images/ID). The bottom row corresponds to results where no filtering was applied.
@@ 5 Conclusions
In this paper, we propose VariFace, a two-stage, diffusion-based pipeline for generating fair and diverse synthetic face datasets for training FR models. We introduce Face Recognition Consistency to refine demographic labels, Face Vendi Score Guidance to improve interclass diversity, and Divergence Score Conditioning to balance the ID preservation-intraclass diversity trade-off. When controlling for the dataset size and the number of images per ID, VariFace consistently outperforms previous synthetic dataset methods across six evaluation datasets and achieves FR performance comparable to real data (Table 1, 2). Furthermore, by scaling synthetic dataset size, VariFace outperforms the real dataset, as well as previous synthetic methods at all dataset sizes (Table 3, 4). Importantly, VariFace addresses fairness concerns, achieving better representation of minority demographic classes demonstrated both through qualitative visualizations (Fig. 4) and quantitative evaluation (Table 2).
@@
With a gap between the FR performances from training on synthetic and real datasets, previous methods have emphasized the benefit of using synthetic data to augment real datasets (Qiu et al., 2021; Bae et al., 2023; Kim et al., 2024). In contrast, we demonstrate for the first time that state-of-the-art FR performance can be achieved when training only on synthetic data. Therefore, our results establish synthetic face datasets as a viable solution for achieving accurate and fair FR performance.
Broader Impact Statement
Over the past decade, there has been significant improvements in automated face recognition performance, with models achieving human-level performance (Taigman et al., 2014; Phillips et al., 2018). The performance improvement is not only due to architectural developments, but perhaps equally, if not more importantly, due to the significant increase in data used to train these models. However, these massive face datasets were obtained by web-scraping and do not have consent from individuals to use their faces for this purpose. To avoid the need to use these datasets, there has been growing interest to develop synthetic face datasets as an alternative. However, the best current synthetic methods, including our proposed method, still rely on real data to train the synthetic models. While this is a better option than directly training on real data, synthetic methods based on CG data or small-scale consented real face data are promising research directions that could eliminate the reliance on massive web-scraped datasets.
Acknowledgments
@@
We are grateful to Prof. Satoshi Hara of the University of Electro-Communications for reviewing this manuscript.
Appendix A CLIP Demographic labeling
A.1 Framework
@@
We adapt the CLIP-IQA framework (Wang et al., 2023) to obtain demographic labels without relying on supervised models (Fig. S1). Unlike CLIP-IQA, we input both the image and its horizontally flipped version into the CLIP image encoder to obtain their corresponding image embeddings. We use the horizontally flipped image as a form of test-time augmentation to improve the robustness of the predictions. For the text encoder, we pass a set of prompts with the structure “A photo of a * face" where * is replaced with [“Male", “Female"], [“Young", “Old"], [“Caucasian", “Asian", “Indian", “African"] for gender, age and race labeling, respectively. Given a set of text embeddings $T=\{t_{1},t_{2},...,t_{n}\}$ and the embedding of an image and its horizontally flipped version $X=\{x_{1},x_{2}\}$, first the cosine similarity is computed:
@@
A.2 Evaluation
To evaluate the benefit of using CLIP compared to supervised models, we compare CLIP with DeepFace (Serengil & Ozpinar, 2021), a library that contains supervised models for race prediction of face images. Furthermore, we evaluate the benefit of including Face Recognition Consistency (FRC) with CLIP predictions.
For evaluation, we use the RFW dataset, which contains 48,000 images evenly divided across the following races: [“Caucasian", “Asian", “Indian", “African"]. DeepFace includes “Latino Hispanic" and “Middle Eastern" as predicted classes, and we relabel these as “Indian" predictions to conform to the RFW race categories. For FRC, we set $K=10$ to account for the small dataset size. The accuracy of the predictions using DeepFace, CLIP, and CLIP-FRC for each race are shown in Fig. S2.